# Emergent Mixture-of-Experts: Can Dense Pre-trained Transformers Benefit from Emergent Modular Structures?

### v1

@@ A.2.2 GLUE
A.3 Computation Cost and Memory Usage
Theoretically, EMoE does not introduce additional parameters compared to its dense counterpart. Although it adds computation in the gating portion within the MoEs layer, it omits a substantial amount of computation within the FFNs layer. For instance, the computation in the gating portion is of the order of $h\times N$, where h represents the model’s hidden size, and N is the number of experts. In contrast, the complete computation in the FFNs layer is of the order of $(h\times h\times 4h)\times 2$, and sparse activations can reduce more than a quarter of this computation. Since $N\ll h$, theoretically, using EMoE within a single block should accelerate the forward pass of the model. However, in real deployment, we have observed that the hardware implementation of MoEs can result in EMoE being, on average, approximately 10% slower than the dense model. Additionally, the memory usage is also slightly higher, by less than 5%, compared to the dense model (16295MB v.s. 15865MB on CoLA, LoRA tuning GPT2-XL).
Each individual experiment was conducted on a single NVIDIA 40G A100 GPU. The training times for different tasks ranged from just over ten minutes (RTE) to more than ten hours (QQP).
A.4 Tabular Results
A.4.1 Full Tables in Full Fine-tuning with Standard Deviation
In this section, we present the mean and variance of experiments conducted with three different random seeds. The Domainbed results are demonstrated in Table 8, full fine-tuning results are in Table 9, LoRA tuning results are in Table 10.
Table 8: Overall out-of-domain performances with different selection criteria. All the reported results are obtained from three independent experiments conducted following the Domainbed codebase. The best result is highlighted in bold.In cases where results are the same, the best result is determined by the smallest standard deviation. EMoE demonstrates comparable results to GMoE.
@@ A.4.1 Full Tables in Full Fine-tuning with Standard Deviation
A.4.2 Full Tables and Figure Data Sources in Analysis
In the Analysis section 5.1, for the sake of clarity, we have transformed tabular data into graphs or retained only a subset of the results. The original and complete results corresponding to them are presented in this section. Figure 3 is summarized from Table 11 and Table 12. The OOD results in Table 3 are from 10. The ablation studies are from Table 14 and Table 15.
Table 11: ID and OOD results of BERT-L for different settings. ”LoRA-to-EMoE” refers to converting a model tuned using standard LoRA into EMoE for testing. On the other hand, ”EMoE-to-LoRA” involves merging a tuned EMoE model back into a standard dense model during testing.
@@ A.4.2 Full Tables and Figure Data Sources in Analysis
Table 12: ID results of BERT-L for different settings. ”Cluster-top” refers to EMoE utilizing $\operatorname{avg-k}$ gating. ”Cluster-not-top” represents a scenario where, during gating, the top-k experts are removed. Similarly, ”Cluster-bottom” involves selecting the bottom-k experts with the lowest scores during gating. ”Random” denotes the approach of randomly selecting key values to construct experts. The terms ”top,” ”not-top,” and ”bottom” have the same meanings as in the cluster situations.
@@
Table 15: Comparison of EMoE and Dense Results with Different Training Data Fraction
@@
Fraction EMoE Dense EMoE Dense EMoE Dense EMoE Dense EMoE Dense EMoE Dense Diff
@@
A.5 More Visualization Results
In this part, we demonstrate more gating visualization results on SST-2, STSB, MRPC, and RTE in Figure 8. These results are consistent with earlier findings: (1) Both $\operatorname{avg-k}$ gating and learned gating converge, as indicated by the lower halves of the plots. (2) $\operatorname{avg-k}$ gating is more stable than learned gating. This could mitigate data inefficiency resulting from inconsistencies in gating across different stages of training (Zuo et al., 2022).
Figure 8: Expert selections during training with distinct gating functions ($\operatorname{avg-k}$ vs learned gate) and expert types (modules from dense vs repetitions of dense). The vertical axis illustrates training steps, with top-down arrangement signifying begin-end; the horizontal axis represents expert selection frequency within 1K steps. (a), (b) and (c) correspond respectively to EMoE, EMoE-learn, and GMoE configurations. The subplots from top to bottom are results for SST-2, STS-B, MRPC, and RTE.

### v2

@@ D.1 GLUE
Appendix E Computation Cost and Memory Usage
E.1 Experiments with public MoEs library
Theoretically, EMoE does not introduce additional parameters compared to the standard model. Although it adds computation in the gating portion within the MoEs layer, it omits a substantial amount of computation within the FFNs layer. For instance, the computation in the gating portion is of the order of $h\times N$, where h represents the model’s hidden size, and N is the number of experts. In contrast, the complete computation in the FFNs layer is of the order of $(h\times h\times 4h)\times 2$, and sparse activations can reduce more than a quarter of this computation. Since $N\ll h$, theoretically, using EMoE within a single block should accelerate the forward pass of the model. However, in real deployment, we have observed that the hardware implementation of MoEs can result in EMoE being, on average, approximately 10% slower than the standard model. The memory usage is also slightly higher, by less than 5%
Each experiment was conducted on a single NVIDIA 40G A100 GPU. The training times for different tasks ranged from just over ten minutes (RTE) to more than ten hours (QQP).
E.2 Experiments with self-implementation
We further find that the increasing wall time and the GPU memory usage come from the public library tutel we used to implement EMoE. We reimplemented our method and observed that EMoE does not require significant additional run time and memory usage. Specifically, we introduce an alternative implementation approach in EMoE where hidden states are used to calculate gate scores after computing the first layer. These scores mask the outputs of unselected experts, mimicking the effect of MoEs. Though this theoretically increases FLOPS compared to traditional MoEs, in practice, the speed is comparable to standard models, as demonstrated in Tab. 72 All experiments about LLMs are conducted on eight NVIDIA 80G A100 GPU.
Appendix F Tabular Results
F.1 Full Tables in Full Fine-tuning with Standard Deviation
This section presents the mean and variance of experiments conducted with three different random seeds. The Domainbed results are demonstrated in Table 13, full fine-tuning results are in Table 14, LoRA tuning results are in Table 15.
Table 13: Overall out-of-domain performances with different selection criteria. All the reported results are obtained from three independent experiments following the Domainbed codebase. The best result is highlighted in bold.In cases where results are the same, the smallest standard deviation determines the best result. EMoE demonstrates comparable results to GMoE.
@@ F.1 Full Tables in Full Fine-tuning with Standard Deviation
F.2 Full Tables and Figure Data Sources in Analysis
In the Analysis section 4, we have transformed tabular data into graphs or retained only a subset of the results for clarity. The original and complete results corresponding to them are presented in this section. Fig. 3 is summarized from Tab. 16 and Tab. 17. The OOD results in Table 1 are from 15. The ablation studies are from Tab. 19, and Tab.20.
Table 16: ID and OOD results of BERT-L for different settings. "LoRA-to-EMoE" refers to converting a model tuned using standard LoRA into EMoE for testing. On the other hand, "EMoE-to-LoRA" involves merging a tuned EMoE model back into a standard standard model during testing.
@@ F.2 Full Tables and Figure Data Sources in Analysis
Table 17: ID results of BERT-L for different settings. "Cluster-top" refers to EMoE utilizing $\operatorname{avg-k}$ gating. "Cluster-not-top" represents a scenario where, during gating, the top-k experts are removed. Similarly, "Cluster-bottom" involves selecting the bottom-k experts with the lowest scores during gating. "Random" denotes the approach of randomly selecting key values to construct experts. The terms "top," "not-top," and "bottom" have the same meanings as in the cluster situations.
@@
Table 20: Comparison of EMoE and Standard Results with Different Training Data Fraction
@@
Fraction EMoE Standard EMoE Standard EMoE Standard EMoE Standard EMoE Standard EMoE Standard Diff
@@
Table 21: Results of top and bottom selection strategies on SST2 and CoLA datasets with different activation ratios. The activation ratio is determined by calculating the proportion of activated neurons that belong to the selected expert among all activated neurons. Meanwhile, the weighted activation ratio is computed by taking the ratio of the sum of activation scores in the selected experts to the sum of the activation scores across the entire FFNs.
Top selection Bottom selection
Dataset Activation Ratio EMoE EMoE-learn EMoE EMoE-learn
SST2
Activation Ratio 32 0.6879 0.6796 0.3121 0.0552
16 0.4317 0.4247 0.124 0.1293
8 0.2616 0.2558 0.0528 0.3218
Weighted Activation Ratio 32 0.731 0.7234 0.269 0.2791
16 0.4953 0.4888 0.1041 0.109
8 0.3304 0.3237 0.0442 0.046
CoLA
Activation Ratio 32 0.6815 0.6788 0.3185 0.3212
16 0.4292 0.4239 0.1295 0.131
8 0.264 0.2582 0.0552 0.0637
Weighted Activation Ratio 32 0.7341 0.4958 0.2659 0.2744
16 0.512 0.3389 0.0468 0.1109
8 0.3589 0.7256 0.1063 0.0488
Table 22: Comparing EMoE and pruning at different states. T-P means training EMoE, pruning experts with lower selection frequency, and evaluating the pruned model. P-T means pruning lower selection frequency experts of an untrained model, then training and evaluating the pruned model.
CoLA SST2
LoRA 63.40 93.16
EMoE 65.33 93.54
Remained Expert T-P P-T T-P P-T
64 65.26 63.4 93.42 93.16
32 65.33 63.52 93.54 93.34
16 65.33 63.42 93.45 93.27
8 64.33 63.42 93.45 93.27
4 63.80 63.21 93.34 93.21
Appendix G More Visualization Results
In this part, we demonstrate more gating visualization results on SST-2, STSB, MRPC, and RTE in Figure 8. These results are consistent with earlier findings: (1) Both $\operatorname{avg-k}$ gating and learned gating converge, as indicated by the lower halves of the plots. (2) $\operatorname{avg-k}$ gating is more stable than learned gating. This could mitigate data inefficiency resulting from inconsistencies in gating across different stages of training (Zuo et al., 2022a).
Figure 8: Expert selections during training with distinct gating functions ($\operatorname{avg-k}$ vs learned gate) and expert types (modules from the standard vs repetitions of the standard). The vertical axis illustrates training steps, with top-down arrangement signifying begin-end; the horizontal axis represents expert selection frequency within 1K steps. (a), (b) and (c) correspond respectively to EMoE, EMoE-learn, and GMoE configurations. The subplots from top to bottom are results for SST-2, STS-B, MRPC, and RTE.